OpenStudy (anonymous):

A solid gold cube with edge of 2" is melted down and recast as a cone with a 3" radius base. Find the height of the cone.

OpenStudy (unklerhaukus):

\[V_{\text{cube}}=s^3\]\[V_{\text{pyramid}}=\frac{bh}3\]a cone is a pyramid with a circular base\[b_{\text{cone}}=\pi r^2\] \[V_{\text{Au}}=V_{\text{cube}}=V_{\text{cone}}\]

OpenStudy (unklerhaukus):

can you see what to do?

OpenStudy (anonymous):

erm i dunno 6.9 duuurrr halp

OpenStudy (unklerhaukus):

first plug in the side length of the cube into V_cube find the volume,

OpenStudy (unklerhaukus):

then solve the volume of the cone for h

OpenStudy (unklerhaukus):

and substitute in the radius r

OpenStudy (anonymous):

and then the answer is what?

OpenStudy (unklerhaukus):

i hav't worked it out, that is your job

OpenStudy (anonymous):

haven't*

OpenStudy (anonymous):

\[v_{cube}=v_{cone}\] \[s^{3}= \frac{ 1 }{ 3 }{\pi}r^{2}h\] \[2^{3}=\frac{ 1 }{ 3 }{\pi}3^{2}h\] \[h=\frac{ 8 }{ 3\pi }\]

OpenStudy (anonymous):

thank you